OpenStudy (anonymous):

On the graph of the equation 5x – 2y = 20, what is the value of the y-intercept?

OpenStudy (mimi_x3):

y-intercept: x=0 5(0) - 2y = 20 -2y = 20 -y = 10 y = -10